# Dynamics of Streamers and Pseudostreamers and Implications for the Solar Wind

### v1

@@ 5 Synthetic In-Situ observation at streamers
Figure 6: In-situ measurement of the electron number density from the Solar wind model: In panel (a), in situ electron number density is sampled over a line segment of length $0.1~R_{\odot}$ above the Helmet streamer stalk at $r\approx 16~R_{\odot}$ (as shown by the white arrow annotation in Fig. 4(a)). The blue line plot shows the evolution of the average number density over sampled points and the background area represents the minimum and maximum value. In panel (b), we follow the same procedure for a $0.1~R_{\odot}$ arc near the Pseudostreamer stalk (as shown in Fig. 5(a)), located at $r\approx 2.5~R_{\odot}$.
@@
The hot, dense plasma released from the closed field regions by the reconnection processes described in the previous section should be transported out into the heliosphere, leading to fluctuations in observable quantities. From our model, we calculate synthetic in situ measurements of the electron number density ($n$) above the stalk of both streamers. This is done by simply calculating $n$ at two sets of points covering the stalks. The in-situ observation points for the Helmet streamer are selected along a $0.1~R_{\odot}$ straight line, positioned at $r\approx 16~R_{\odot}$ (see the white arrow marker in Figure 4(a)). We chose this location because Parker Solar Probe recently measured in-situ plasma quantities at a similar distance from the Sun. In Figure 6(a), the variation of the electron number density at this location is shown. Both the spatially averaged profile (blue-solid curve) and the minimum to maximum number density range (blue shaded region) show a clear periodicity of 12.07 hours with varying amplitudes. This density oscillation over the Helmet streamer stalk is due to the quasi-periodic plasmoid ejections (see Section 4), which propagate away from the Sun and across the measurement points (Figure 4). During its 8th solar encounterParker Solar Probe was located at $r\approx 16~R_{\odot}$ from the Sun, and using quasi-thermal noise spectroscopy, the local electron density was measured at around  $6\times 10^{3}-7\times 10^{3}$ cm-3 (kasper2021; Kruparova2023). This result matches in order of magnitude with our simulated values at a similar distance. To measure the open-field density variation induced by the Pseudostreamer dynamics, we select in-situ observation points along an arc length of $0.1~R_{\odot}$, positioned above the Pseudostreamer stalk at $r\approx 2.5~R_{\odot}$ (see the white arrow marker in Figure. 5(a)). The averaged electron density over the selected points (dashed magenta line in Figure. 6(b)) exhibits a fluctuating nature with time. The maximum and minimum values of the electron density for the same selected points (magenta shaded region of Figure 6(b) also show a large spread in the density distribution. This fluctuating temporal behaviour of the electron density profile can be described in terms of the induced IMR process at the Pseudostreamer. As we discuss in Subsection 4.2, dense plasma elements from the closed magnetic arcades of the Pseudostreamer can jump onto open field lines through the IMR mechanism. Then these dense plasma parcels propagate outward as plasmoids with the solar wind and contribute to a bursty electron density measurement. In the latest in-situ observations of PSP, fluctuating plasma characteristics have been well observed for several close encounters with the sun. Although the origin of the fluctuations is not completely understood, we believe several observed aspects can be described qualitatively by simple IMR models like our solar wind model.
6 Discussion
@@
In this paper our goal was to explore the interaction of closed and open magnetic field (and associated plasma) under generic conditions on the Sun. That is, we are interested not in the rare, large eruptions from the closed field, but rather in the quiescent behaviour driven by the continual sub-surface convective flows. In order to extend previous studies of interchange reconnection – and in particular to enable closer connections to be made to the observations – we have included in our model a chromospheric layer, as well as field-aligned thermal conduction, radiative losses, and empirical heating functions. With these physical effects included, we obtain a quasi-steady state with an outflowing wind in which the plasma in the closed-field region is both hotter (by a range of 1.40–1.72) and denser (by a factor of $\approx$ 8) than in the adjacent open field, broadly consistent with observations. As such, when magnetic flux changes connectivity from closed to open during the simulations, the hot, dense plasma can expand into the cooler, more tenuous surroundings along the newly open flux tube.
@@
Even in the absence of any boundary driving, both the helmet streamer and pseudostreamer show an ongoing, periodic dynamics. For the helmet streamer this is expected based on numerous previous studies, but for the pseudostreamer we speculate that the dynamics are generated/maintained by the plasma thermodynamics in the closed field. As such, it will be of interest in future to explore whether this behaviour survives when different heating functions are employed, for different sizes of pseudostreamer, in three dimensions, etc. In both cases, the dynamics become more pronounced when boundary shearing is introduced to mimic the effects of supergranular driving.
@@
At the helmet streamer, the interaction between closed and open flux takes the form of a “breathing” cycle in which the closed flux contracts and expands, and plasmoids are released along the heliospheric current sheet. The period of this “pinching off” of plasmoids in our simulations is around 13 hours, and leads to a similar periodicity in density fluctuations in the vicinity of the heliospheric current sheet. However, we note that a preliminary exploration indicates that the periodicity depends on the grid refinement that we permit along the heliospheric current sheet. This is in spite of the fact that the simulation never reaches maximum refinement in that region (when the maximum refinement level exceeds 5). Therefore, in a future study, we will make a systematic study where we vary the grid refinement level and strategy to seek a limiting value for this period.
@@
By contrast to the helmet streamer, at the pseudostreamer, the plasma exchange between open and closed field must be mediated by interchange reconnection in which, in any finite time window, an equal amount of magnetic flux is opening on one side and closing on the other side of the pseudostreamer (no net opening or closing). The pseudostreamer exhibits a back-and-forth motion, driving interchange reconnection alternately on its opposite flanks. The resulting release of hot, dense plasma leads to density fluctuations in the open field that are significantly larger above the helmet streamer due to the persistence of the plasmoids there.
@@
Our model demonstrates that plasma is continually being released into the heliosphere from both streamers and pseudostreamers. As a result, across the S-Web we should expect to see plasma with mixed properties in terms of the magnetic connectivity at their origins in the low corona. Fluctuations in plasma properties above helmet streamers should by larger in magnitude, because plasmoids that form during the reconnection process can be carried out into the heliosphere intact. Even in 3D – where these plasmoids are not strictly magnetically closed – the magnetic flux rope structure of the plasmoid will at least inhibit mixing between the newly released plasma and the ambient wind. By contrast, plasma released from the pseudostreamer is not magnetically confined, since due to the field orientation, the (much smaller) plasmoids formed during interchange reconnection there are quickly eroded by reconnection between the plasmoid field and ambient open field. As such, the plasma released from the closed field region beneath the pseudostreamer quickly mixes with the adjacent open-field plasma.
@@
There are a number of limitations to the present study, that we plan to overcome in future works. (i) The most obvious is the 2.5D nature of the simulations. 3D offers a much broader range of magnetic – and therefore plasma – dynamics, particularly in terms of the formation of new, open flux (e.g., aslanyan2022). In addition, as mentioned above, plasmoids are not magnetically closed structures in 3D, with important implications for plasma transport to larger radii. (ii) Our solar wind is rather slow compared to the observed value: faster winds can be obtained by, for example, adding an additional Alfvén wave term to the momentum equation. (iii) As mentioned above, with the present grid refinement strategy we could not find a definitive, resolution-independent periodicity for the fluctuations induced. In future we will explore refining on, for example, the current density. (iv) Tracking plasma parcels released from the closed field into the open field would permit a more detailed comparison with observations, and will also be implemented in the future.
@@
S.D. and D.P. gratefully acknowledge support through an Australian Research Council Discovery Project (DP210100709). The authors acknowledge the helpful discussions with Craig Johnston and Cooper Downs. S.D. is grateful to Chun Xia, Yang Guo, and Rony Keppens for extensive discussions and suggestions regarding various numerical and physical aspects of the solar wind modelling.

### v2

@@ V Synthetic In-Situ observation at streamers
Figure 6: In-situ measurement of the electron number density from the Solar wind model: In panel (a), in situ electron number density is sampled over a line segment of length $0.1~R_{\odot}$ above the Helmet streamer stalk at $r\approx 16~R_{\odot}$ (as shown by the white arrow annotation in Fig. 4(a)). The cyan line plot shows the evolution of the average number density over sampled points and the background area represents the minimum and maximum value of density over the line segment. In panel (b), we follow the same procedure for the equal angular width arc as panel (a), which is $0.01~R_{\odot}$ line segment at $r\approx 2.5~R_{\odot}$, located near the Pseudostreamer stalk (see the arrow in Fig. 5(a))
@@
The hot, dense plasma released from the closed field regions by the reconnection processes described in the previous section should be transported out into the heliosphere, leading to fluctuations in observable quantities. From our model, we calculate synthetic in situ measurements of the electron number density ($n$) above the stalk of both streamers. This is done by simply calculating $n$ at two sets of points covering the stalks. The in-situ observation points for the Helmet streamer are selected along a $0.1~R_{\odot}$ straight line, positioned at $r\approx 16~R_{\odot}$ (see the white arrow marker in Figure 4(a)). We chose this location because Parker Solar Probe recently measured in-situ plasma quantities at a similar distance from the Sun. In Figure 6(a), the variation of the electron number density at this location is shown. Both the spatially averaged profile (blue-solid curve) and the minimum to maximum number density range (blue shaded region) show a clear periodicity between 11.5–14.0 hours with varying amplitudes. This density oscillation over the Helmet streamer stalk is due to the quasi-periodic plasmoid ejections (see Section IV), which propagate away from the Sun and across the measurement points (Figure 4). During its 8th solar encounter, Parker Solar Probe was located at $r\approx 16~R_{\odot}$ from the Sun, and using quasi-thermal noise spectroscopy, the local electron density was measured at around  $6\times 10^{3}-7\times 10^{3}$ cm-3 (Kasper et al., 2021; Kruparova et al., 2023). This result matches in order of magnitude with our simulated values at a similar distance.
To measure the open-field density variation induced by the Pseudostreamer dynamics, we select in-situ observation points along an equal angular width arc as the Helmet streamer case above the Pseudostreamer stalk, which is $0.01~R_{\odot}$ arc length at $r\approx 2.5~R_{\odot}$ (see the white arrow marker in Figure. 5(a)). The averaged electron density over the selected points (dashed magenta line in Figure. 6(b)) exhibits a fluctuating nature with time. This fluctuating temporal behaviour of the electron density profile can be described in terms of the induced IMR process at the Pseudostreamer. As we discuss in Subsection IV.2, dense plasma elements from the closed magnetic arcades of the Pseudostreamer can jump onto open field lines through the IMR mechanism. Then these dense plasma parcels propagate outward as plasmoids with the solar wind and contribute to a bursty electron density measurement. In the latest in-situ observations of PSP, fluctuating plasma characteristics have been well observed for several close encounters with the sun. Although the origin of the fluctuations is not completely understood, we believe several observed aspects can be described qualitatively by simple IMR models like our solar wind model.
Figure 7: A temporal map of the synthetic electron Strahl: it is computed at a point located above the Helmet streamer stalk at $r=16~R_{\odot}$. Here, $\Phi$ represents the relative angle between the magnetic field and the heat flux carried by the strahl. The RGB vertical segment corresponds to heat fluxes from the closed-magnetic field lines. A blue vertical patch indicates the absence of strahl, which is associated with the plasmoids. The narrow region between the black vertical lines represents the fluxes that propagate away from the sun.
To gain further insights in terms of the in-situ observables, we investigate a synthetic Strahl map from our model. Strahl is a measure of suprathermal electron heat flux that indicates the magnetic field orientation and the nature of magnetic connectivity (Feldman et al., 1975; Pilipp et al., 1987). It is routinely observed in situ by several space-based observation facilities, e.g., Wind, Helios, Parker Solar Probe, Solar Orbiter, etc. From our model, we compute the synthetic in-situ strahl map (Fig. 7) for a fixed point above the Helmet streamer stalk at $r=16~R_{\odot}$ by evaluating the magnetic field line connectivity for the field line passing through that point, as a function of time.
The map clearly shows the source of strahl electron from three different types of magnetic structures; (i) strahl from closed magnetic loop (denoted by RGB vertical sections) (ii) no strahl from magnetically isolated plasmoids (blue vertical strip) (iii) strahl from open flux, where the foot point is anchored at the negative polarity region of the chromosphere (narrow area between vertical black lines). In the strahl map, we find repeated occurrences of no strahl phases, due to the quasi-periodic plasmoid ejection at the HCS. Since the point we have selected for this calculation is to one side of the HCS, we note that when the field is open, it is always connected to negative polarity at the base (corresponding to a strahl signal around $\Phi=180$). We emphasise that our simulation lacks the complex physics to generate suprathermal electrons – the map is a visualisation of the field line connectivity alone, but it reveals the periodic strahl variations that we would expect to find in observations for such a field evolution.
To explore the nature of the propagating disturbances in our simulations, we compute the Alfvénicity or normalized cross-helicity ($\sigma$). This quantity measures the correlation between velocity and magnetic field fluctuations following Alfvén wave properties (Barnes, 1979; Matthaeus and Goldstein, 1982; Tu and Marsch, 1995). The definition of Alfvénicity is as follows,
$\sigma=<2{\boldsymbol{\delta}}\mbox{$u$}{}(t)\cdot{\boldsymbol{\delta}}v_{A}(t)>/<(\boldsymbol{\delta}\mbox{$u$}{}(t)^{2}+\boldsymbol{\delta}v_{A}(t)^{2})>$, where $\boldsymbol{\delta}\mbox{$u$}{}(t)=\mbox{$u$}{}(t)-<\mbox{$u$}{}(t)>$,
$\boldsymbol{\delta}v_{A}(t)=(\mbox{$B$}{}(t)-<\mbox{$B$}{}(t)>)/\sqrt{4\pi<\rho(t)>}$ in cgs unit and <><~> represents a temporal average, chosen here to be over half an hour period. We calculate fluctuations in the velocity field $(\boldsymbol{\delta}\mbox{$u$}{})$ and magnetic field $(\boldsymbol{\delta}\mbox{$B$}{})$ with 1-minute cadence. In Figure 8, the Alfvénicity is shown at $r=16~R_{\odot}$ over a range of colatitudes, covering both the Helmet and Pseudostreamer. To remove any artifact of the imposed polar boundary conditions, we neglect regions close to the poles. We see that the Alfvénicity changes sign across the HCS, in general taking negative (positive) values when the temporally averaged radial magnetic field is positive (negative). These characteristics affirm that all of the fluctuations embedded in the solar wind are propagating away from the sun at $16~R_{\odot}$. A very high level of Alfvénicity $\sigma=-0.9$ is detected in the colatitude range of 110-120 degrees (Figure 8); this location corresponds to the Pseudostreamer in the lower atmosphere. This is one of the crucial results from our model that suggests Pseudostreamers can contribute to a strong level of Alfvénic tubulence by exciting Alfvénic waves during the IMR process. Recent observations from the Solar Orbiter reveal that a Pseudostreamer does excite a high level of Alfvénicity (D’Amicis et al., 2025) in the slow solar wind. It is also intriguing that fluctuations that originate near the apex of the Pseutostreamer at $r\approx 2.2R_{\odot}$ in our model can survive to a much longer distance at $16~R_{\odot}$ and show their signatures in terms of Alfvénicity measurement. Unlike the Pseudostreamer, the Helmet streamer does not exhibit any sign of high Alfvénic activity. Finally, it should be noted that the fine features of the Alfvénicity curve strongly depend on the dynamic background state and the time period over which fluctuations are computed, but the general nature of the curve remains the same.
Figure 8: Variation of Alfvénicity measurement $(\sigma)$ as a function of colatitude is represented with the black curve at $r=16~R_{\odot}$. The radial component of the magnetic field is averaged over half an hour duration and shown by the blue curve.
VI Discussion
@@
In this paper our goal was to explore the interaction of closed and open magnetic field (and associated plasma) under generic conditions on the Sun. That is, we are interested not in the rare, large eruptions from the closed field, but rather in the quiescent behaviour driven by the continual sub-surface convective flows. In order to extend previous studies of interchange reconnection – and in particular to enable closer connections to be made to the observations – we have included in our model a chromospheric layer, as well as field-aligned thermal conduction, radiative losses, and empirical heating functions. With these physical effects included, we obtain a quasi-steady state with an outflowing wind in which the plasma in the closed-field region is both hotter (by a factor of 1.40–1.72) and denser (by a factor of $\approx$ 8) than in the adjacent open field, broadly consistent with observations. As such, when magnetic flux changes connectivity from closed to open during the simulations, the hot, dense plasma can expand into the cooler, more tenuous surroundings along the newly open flux tube.
@@
Even in the absence of any boundary driving, both the helmet streamer and pseudostreamer show an ongoing, periodic dynamics. For the helmet streamer this is expected based on numerous previous studies, but for the pseudostreamer, we speculate that the dynamics are generated/maintained by the plasma thermodynamics in the closed field, and have thus not been observed before in simulations employing a simplified treatment of the thermodynamics. As such, it will be of interest in future to explore whether this behaviour survives when different heating functions are used, for different sizes of pseudostreamer, in three dimensions, etc. For both the helmet streamer and pseudostreamer, the dynamics become more pronounced when boundary shearing is introduced to mimic the effects of supergranular driving.
@@
At the helmet streamer, the interaction between closed and open flux takes the form of a “breathing” cycle in which the closed flux contracts and expands, and plasmoids are released along the heliospheric current sheet. The period of this “pinching off” of plasmoids in our simulations is around 11.5–14.0 hours, and leads to a similar periodicity in density fluctuations in the vicinity of the heliospheric current sheet. The imprint of these quasi-periodic plasmoid ejections is also well captured in the in-situ diagnostics. e.g., the Strahl map. However, we note that a preliminary exploration indicates that the periodicity depends on the grid refinement that we permit along the heliospheric current sheet. This is in spite of the fact that the simulation never reaches maximum refinement in that region (when the maximum refinement level exceeds 5). Therefore, in a future study, we will make a systematic study where we vary the grid refinement level and strategy to seek a limiting value for this period.
@@
By contrast to the helmet streamer, at the pseudostreamer, the plasma exchange between open and closed field must be mediated by interchange reconnection in which, in any finite time window, an equal amount of magnetic flux is opening on one side and closing on the other side of the pseudostreamer (no net opening or closing). The pseudostreamer exhibits a back-and-forth motion, driving interchange reconnection alternately on its opposite flanks, even though in our simulation the driving is maintained in a constant direction. The resulting release of hot, dense plasma leads to density fluctuations in the open field that are significantly larger above the helmet streamer due to the persistence of the plasmoids there. An important signature of interchange magnetic reconnection that is implied by our simulations is the degree of Alfvénic fluctuations in the solar wind. We predict that a strong level of Alfvénicity/Cross Helicity would be detected over the region of Pseudostreamers, originating from the reconnection site near the apex of the same streamer.
@@
Our model demonstrates that plasma is continually being released into the heliosphere from both streamers and pseudostreamers. As a result, across the S-Web we should expect to see plasma with mixed properties in terms of the magnetic connectivity at their origins in the low corona. Fluctuations in plasma properties above helmet streamers should by larger in magnitude, because plasmoids that form during the reconnection process can be carried out into the heliosphere intact. Even in 3D – where these plasmoids are not strictly magnetically closed – the magnetic flux rope structure of the plasmoid will at least inhibit mixing between the newly released plasma and the ambient wind. In contrast, plasma released from the pseudostreamer is not magnetically confined, since due to the field orientation, the (much smaller) plasmoids formed during interchange reconnection there are quickly eroded by reconnection between the plasmoid field and ambient open field. As such, the plasma released from the closed field region beneath the pseudostreamer quickly mixes with the adjacent open-field plasma. Despite the decay in plasmoid structures, the signatures of the interchange reconnection process are carried to several solar radii distances by the propagating Alfvénic waves.
@@
There are a number of limitations to the present study, that we plan to overcome in future works. (i) The most obvious is the 2.5D nature of the simulations. 3D offers a much broader range of magnetic – and therefore plasma – dynamics, particularly in terms of the formation of new, open flux (e.g., Aslanyan et al., 2022). It may well be that in 3D, more turbulence is generated in situ in the open field rather than primarily at the open-closed boundary. This may affect the level of Alfvénicity measured in the wind. In addition, as mentioned above, plasmoids are not magnetically closed structures in 3D, with important implications for plasma transport to larger radii. Due to this change for the nature of magnetic connectivity in 3D, the strahl from these plasmoids would be categorized as a contribution from the open magnetic field. (ii) Our solar wind is rather slow compared to the observed value: faster winds can be obtained by, for example, adding an additional Alfvén wave term to the momentum equation. (iii) As mentioned above, with the present grid refinement strategy we could not find a definitive, resolution-independent periodicity for the fluctuations induced. In future we will explore refining on, for example, the current density. (iv) Tracking plasma parcels released from the closed field into the open field would permit a more detailed comparison with observations, and will also be implemented in the future.
@@
S.D. and D.P. gratefully acknowledge support through an Australian Research Council Discovery Project (DP210100709). This research was supported by the Australian Government’s National Collaborative Research Infrastructure Strategy (NCRIS), with access to computational resources provided by the National Computational Infrastructure (NCI) through the National Computational Merit Allocation Scheme (NCMAS). S.D. and D.P. also acknowledge the generous computing grant provided by the AAL Supercomputer Time Allocation Committee (ASTAC). SKA acknowledges support from a NASA LWS grant to the University of Michigan. The authors acknowledge the helpful discussions with Craig Johnston and Cooper Downs. S.D. is grateful to Chun Xia, Yang Guo, and Rony Keppens for extensive discussions and suggestions regarding various numerical and physical aspects of the solar wind modelling.